OpenStudy (anonymous):

The average (mean) value of tan x on the interval from x = 0 to x = pi/3 is

OpenStudy (anonymous):

I thought the mean value theorem was f'(c) = f(b)-f(a)/b-a I plugged in the values and I got \[f'(c) = \sqrt{3}/(\pi/3)\]

OpenStudy (anonymous):

But it wasn't one of the answer choices. Was I supposed to do something else?

hartnn (hartnn):

what are you answer choices ?

hartnn (hartnn):

\(3 \sqrt 3/ \pi\) ?

OpenStudy (anonymous):

a. ln(1/2) b. (3/pi)ln2 c. ln2 d. rad(3) / 2 e. 9/pi

hartnn (hartnn):

why are there so many ln 2 s is the function 'tan x' only ?

OpenStudy (anonymous):

Yes

hartnn (hartnn):

i kinda hope it was tan^2 x

OpenStudy (anonymous):

I don't know. I wrote it exactly how my book had it. And I don't think the book would make too many errors

hartnn (hartnn):

lets see what parth has to say :)

Parth (parthkohli):

Pardon me, but isn't the mean value of a function given by\[\dfrac{\int_{a}^{b} f(x) dx}{b-a}\]Then,\[\frac{-\ln |\cos \pi/3| + \ln |\cos 0|}{\pi /3}\]\[= \frac{-\ln |1/2|}{\pi /3}\]\[= 3\ln (2)/\pi \]

OpenStudy (anonymous):

Ya now I'm pretty sure thats what I should've done

Parth (parthkohli):

Because AFAIK, the right side of the mean value theorem is not actually the mean value of the function.

OpenStudy (anonymous):

I just saw the word mean and applied the mean value theorem. My bad >.<

OpenStudy (nincompoop):

you're mean apply MVT

hartnn (hartnn):

i was thinking something else \(\LARGE \dfrac{f(b)-f(a)}{b-a}\)